OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 30 miles per gallon of gas and the second has a fuel efficiency of 40 miles per gallon of gas. During one particular week, the two cars went a combined total of 1700 miles, for a total gas consumption of 50 gallons. How many gallons were consumed by each of the two cars that week

OpenStudy (anonymous):

@misty1212 I need help :(((( with a few word problems like this :(

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i hate word problems don't you?

OpenStudy (misty1212):

we need a variable, maybe two or we could get away with one, your choice

OpenStudy (anonymous):

ugh yes I hate this dislike them basically I just don't like them in general :( @misty1212

OpenStudy (dan815):

i love them :)

OpenStudy (misty1212):

ok say \(x\) is the distance travelled by the 30 mph car and y the distance traveled by the 40 mph car then we can make two equations \[x+y=1700\] is the easy one

OpenStudy (misty1212):

the amount of gas used by the 30 mph car is \(\frac{x}{30}\) and the amount used by the 40 mph car is \(\frac{y}{40}\) so the next equation is \[\frac{x}{30}+\frac{y}{40}=50\]

OpenStudy (misty1212):

you know how to solve that system? \[x+y=1700\\ \frac{x}{30}+\frac{y}{40}=50\]